1)A steel wire ABC 16m long having cross sectional area of 4mm2 weighs 20N . If the modulus of elasticity for the wire material is 200Gpa find the deflections at C and B

2)A max stress in a 6.36mm steel wide flange beam 100mm long moment of inertia 362mm4,modulus of elasticity 29,000Gpa with a load of 15kN.Calculate the max stress and deflection
1
Expert's answer
2019-07-08T05:28:56-0400

1) Deflection at C due to self weight of wire AC = dlc

"dl_c = \frac { Wl} {2AE} (1)"

Using (1) we get: dlc= 0.2 mm


Deflection at B:

Now deflection at B is due to two reasons: i) due to self

weight of AB and ii) due to weight of BC.


"dl_B = \frac { 0.5W×0.5l} {2AE} + \frac { 0.5W×0.5l} {AE} (2)"

Using (2) we get: dlB= 0.15 mm


2)

Maximum strees is given by formula

"σ= \frac { y q L^2} {8I} (1)"

where y is distance to extreme point from neutral, q is the uniform load per length unit of beam (N/m, N/mm, lb/in), L is the length of beam (m, mm, in), I is the moment of inertia of the cross section


"q=\frac { W} {L} (2)"

In our case, y=6.36mm, L=100 mm, W=15kN, I=362mm4

Using (1) we get: σ = 4.95×104 Pa


Maximum deflection is given by formula

"δ= \frac { 5 q L^4} {384 E I} (3)"

In our case, E=29,000Gpa

Using (3) we get: δ = 6×10-5 mm